How much water vapour exist in a 105-m³ room on a day when the relative humidity in the room is 32 percent and the room temperature is 20.0°C ? Saturated air at 20°C contains 17.12 g/m³ of water.
1
Expert's answer
2020-11-20T09:30:53-0500
"\rho=0.32(17.12)=5.48\frac{g}{m^3}"

"m=\rho V=(105)(5.48)=575\ g"